# IQ-Aware Precoding for Atomic MIMO Receivers

### v1

@@ IV-A Precoding architecture
whose covariance matrix is $\bar{{\mathbf{Q}}}=\frac{1}{2}{\mathbf{F}}_{r}{\mathbf{F}}_{r}^{T}$. Due to the complex product relationship between ${\mathbf{F}}$ and ${\mathbf{s}}$ that produces ${\mathbf{x}}$, the equivalent linear mapping ${{\mathbf{F}}}_{r}$ that constructs $\bar{{\mathbf{x}}}$ has dependent diagonal and off-diagonal block components. This dependency makes it impossible to assign ${\mathbf{F}}_{r}$ with the principal singular vectors of the real-valued channel $\bar{{\mathbf{H}}}$ in (15). Thereby, the classical complex precoder is not suitable to the transmission model (8).
@@
To address this issue, we propose to replace the conventional complex digital precoding with an IQ-aware digital precoding approach as presented in Fig. 3 (b). Unlike the joint manipulation of baseband symbols using the real and imaginary components of the complex precoder ${\mathbf{F}}$, our approach employs four real-valued matrices, ${\mathbf{F}}_{k\ell}\in\mathbb{R}^{N_{t}\times N_{s}},k\in\{1,2\},\ell\in\{1,2\}$, to independently process ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$. By this means, the signals entering the IQ modulator become ${\mathbf{F}}_{11}{\mathbf{s}}_{I}+{\mathbf{F}}_{12}{\mathbf{s}}_{Q}$ and ${\mathbf{F}}_{21}{\mathbf{s}}_{I}+{\mathbf{F}}_{22}{\mathbf{s}}_{Q}$, respectively. After combining these IQ components, the transmit signal in complex form is given as
${\mathbf{x}}={\mathbf{F}}_{11}{\mathbf{s}}_{I}+{\mathbf{F}}_{12}{\mathbf{s}}_{Q}+j({\mathbf{F}}_{21}{\mathbf{s}}_{I}+{\mathbf{F}}_{22}{\mathbf{s}}_{Q}),$ (26)
while the real form is
$\bar{{\mathbf{x}}}=\left(\begin{array}[]{c}{\mathbf{x}}_{I}\\ {\mathbf{x}}_{Q}\end{array}\right)=\underbrace{\left(\begin{array}[]{cc}{\mathbf{F}}_{11}&{\mathbf{F}}_{12}\\ {\mathbf{F}}_{21}&{\mathbf{F}}_{22}\end{array}\right)}_{\bar{{\mathbf{F}}}}{\left(\begin{array}[]{c}{\mathbf{s}}_{I}\\ {\mathbf{s}}_{Q}\end{array}\right)}.$ (33)
Here, $\bar{{\mathbf{F}}}\in\mathbb{R}^{2N_{t}\times 2N_{s}}$ is named the “IQ-aware digital precoder”. Evidently, the complex signal ${\mathbf{x}}$ in (26) is no longer the product of a complex-valued precoder and complex symbol, like in (18). More importantly, the equivalent linear mapping $\bar{{\mathbf{F}}}$ that generates $\bar{{\mathbf{x}}}$ has independent block components, thereby imposing no additional constraints on the structure of covariance matrix $\bar{{\mathbf{Q}}}=\frac{1}{2}\bar{{\mathbf{F}}}\bar{{\mathbf{F}}}^{T}$. Henceforth, the proposed IQ-aware digital precoding offers more flexible manipulation to the radiated signal ${\mathbf{x}}$ than the complex digital precoding. Therefore, it has the potential to achieve the instantaneous capacity of atomic receivers.
IV-B Precoding method
@@
We now seeks an IQ-aware digital precoder that maximizes the mutual information. Let $\bar{{\mathbf{V}}}\in\mathbb{R}^{2N_{t}\times 2N_{t}}$ denote the right-singular matrix of $\bar{{\mathbf{H}}}$, and let $\{\bar{\lambda}_{1},\cdots,\bar{\lambda}_{\min\{N_{r},2N_{t}\}}\}$ represent the corresponding singular values of $\bar{{\mathbf{H}}}$ in a descending order. According to the standard analysis of MIMO capacity, the optimal IQ-aware digital precoder $\bar{{\mathbf{F}}}$ is given by:
@@
The power allocation matrix $\bar{{\mathbf{P}}}\overset{\Delta}{=}\mathrm{diag}\{\bar{p}_{1},\cdots,\bar{p}_{2N_{s}}\}$ is obtained using the water-filling principle:
$\bar{p}_{k}=\left(\mu-\frac{\sigma^{2}}{\bar{\lambda}_{k}^{2}}\right)^{+},\>k\in\{1,\cdots,2N_{s}\},$ (37)
where the water level $\mu$ is adjusted to satisfy the total power constraint $\mathsf{Tr}\{\bar{{\mathbf{Q}}}\}=\mathsf{Tr}\{\bar{{\mathbf{F}}}\bar{{\mathbf{F}}}^{T}\}=\mathsf{Tr}\{\bar{{\mathbf{V}}}_{:,1:2N_{s}}\bar{{\mathbf{P}}}\bar{{\mathbf{V}}}_{:,1:2N_{s}}^{T}\}=\sum_{k=1}^{2N_{s}}\bar{p}_{k}=P$. Consequently, by processing baseband symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$ using $\{{\mathbf{F}}_{11},{\mathbf{F}}_{12},{\mathbf{F}}_{21},{\mathbf{F}}_{22}\}$, the system achieves its instantaneous capacity.
Last, by substituting $\bar{{\mathbf{F}}}$ into (15), the achieved capacity of atomic MIMO systems, with $N_{s}\leq\min\{N_{r}/2,N_{t}\}$, is expressed as
$C=\sum_{k=1}^{2N_{s}}\frac{1}{2}\log\left(1+\frac{\bar{\lambda}_{k}^{2}\bar{p}_{k}}{\sigma^{2}}\right).$ (38)
Then, we evaluate the asymptotic behavior of $C$. At high SNRs, the water-filling principle tends to allocate equal powers to all data streams, i.e., $\bar{p}_{k}=\frac{P}{2N_{s}}$. Thereafter, the instantaneous capacity is asymptotically
$C=N_{s}\log\frac{\mathrm{SNR}}{N_{s}}+\sum_{k=1}^{2N_{s}}\frac{1}{2}\log\frac{\bar{\lambda}_{k}^{2}}{2},\quad(\text{SNR}\rightarrow\infty).$ (39)
Particularly, when the number of multiplexed data streams approaches the maximum parallel spatial channels, i.e., $N_{s}=\min\{N_{r}/2,N_{t}\}$, the DoF of $C$ is calculated to be $\lim_{\text{SNR}\rightarrow\infty}\frac{C}{\log\text{SNR}}=\min\{N_{r}/2,N_{t}\}$. This DoF is consistent with the analytical result of ergodic capacity, (17), without transmitter side CSI, demonstrating the effectiveness of the derived DoF.
@@ V IQ-Aware Hybrid Precoding
In the preceding section, we have demonstrated that the capacity of atomic MIMO systems can be achieved using IQ-aware digital precoding. However, this solution presents additional challenges. Notably, optimal IQ-aware digital precoding has to manipulate the IQ symbols independently before the IQ modulator. This requirement mandates its implementation on a digital baseband processor, because analog circuits, like phase shifters and power amplifiers, cannot distinguish the combined IQ symbols. Consequently, our IQ-aware digital precoding approach necessitates a fully digital precoding architecture, where the number of RF chains is equal to the number of antennas. However, employing a large number of RF chains results in prohibitive hardware costs and power consumption, especially in massive MIMO systems. To mitigate these challenges, current 5G base stations (BSs) commonly adopt the hybrid analog and digital precoding architectures, presented in Fig. 4. For this architecture, precoding is executed by a low-dimensional digital processor followed by a high-dimensional analog PS network. In this typical scenario, the IQ-aware precoder proposed in (36) is clearly inapplicable due to its reliance on digital processing.
To tackle this challenge, we explore IQ-aware hybrid precoding in this section, incorporating a low-dimensional IQ-aware digital precoder in the baseband processor with an analog PS network in the RF front-end. These components are jointly optimized to approach the optimal IQ-aware digital precoder in (36). As illustrated in Fig. 4, our investigation focuses on two prevalent hybrid precoding architectures: the FC and SC PS networks [34].

### v2

@@ IV-A1 Sub-optimality of traditional digital precoding
whose covariance matrix is $\bar{{\mathbf{Q}}}=\mathsf{E}(\bar{{\mathbf{x}}}\bar{{\mathbf{x}}}^{T})=\frac{1}{2}{\mathbf{F}}_{r}{\mathbf{F}}_{r}^{T}$. The optimality of the precoder structure of ${\mathbf{F}}$ for traditional MIMO and its sub-optimality for atomic MIMO are explained as follows.
Figure 2: (a) Classical complex-valued digital precoding for classical receivers and (b) proposed IQ-aware digital precoding for atomic receivers.
In a traditional MIMO system, the CSCG distribution is capacity achieving. Thus given CSCG distributed data symbols, ${\mathbf{s}}$, as input, a linear complex-valued precoder, ${\mathbf{F}}$, that retains the CSCG distribution can be optimal. Specifically, the optimal ${\mathbf{F}}$ is aligned with the principal singular-vectors of the channel ${\mathbf{H}}$. On the contrary, the CSCG precoded symbols cannot achieve the channel capacity of an atomic MIMO system, making the precoder ${\mathbf{F}}$ sub-optimal. To see the sub-optimality, one can observe from (22) that the real-equivalence of ${\mathbf{F}}$, ${\mathbf{F}}_{r}$, has a block-coupled constraint, i.e., its diagonal block components are identical while its off-diagonal block components are opposite. Obviously, introducing this structural constraint reduces the channel capacity of atomic MIMO in (14). To elaborate, ${\mathbf{F}}_{r}$ under the constraint cannot always align with the principal singular-vectors of the real channel $\bar{{\mathbf{H}}}$ to achieve its capacity. For example, we consider a simple scenario: $N_{r}=N_{t}=2$ and $N_{s}=1$. The $2\times 2$ complex channel matrix is ${\mathbf{H}}=[2,j;j,1]$ and the reference signal is ${\mathbf{r}}=[1;1]$. In this context, the $2\times 4$ real channel matrix is $\bar{{\mathbf{H}}}=[{\mathbf{H}}_{I},-{\mathbf{H}}_{Q}]=[2,0,0,-1;0,1,-1,0]$. The principal singular vectors of $\bar{{\mathbf{H}}}$ are calculated as $\bar{{\mathbf{v}}}_{1}=[-0.8944,0,0,0.4472]^{T}$ and $\bar{{\mathbf{v}}}_{2}=[0,-0.7071,0.7071,0]^{T}$. Evidently, the singular vectors $\bar{{\mathbf{v}}}_{1}$ and $\bar{{\mathbf{v}}}_{2}$ cannot form a precoding matrix satisfying the block-coupled constraint of ${\mathbf{F}}_{r}$. In this context, precoder ${\mathbf{F}}_{r}$ and its complex-equivalence ${\mathbf{F}}$ is not optimal.
IV-A2 Proposed IQ-aware digital precoding
@@
The sub-optimality of conventional precoding motivates the proposed IQ-aware digital precoding in Fig. 2 (b), which relaxes the block-coupled constraint and thereby is capable of generating general-Gaussian symbols despite CSCG input. Specifically, compared with the joint manipulation of baseband IQ symbols using the complex precoder ${\mathbf{F}}$, the IQ-aware precoding employs four real-valued matrices, ${\mathbf{F}}_{k\ell}\in\mathbb{R}^{N_{t}\times N_{s}},k\in\{1,2\},\ell\in\{1,2\}$, to independently process ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$. By this means, the signals entering the IQ modulator become ${\mathbf{F}}_{11}{\mathbf{s}}_{I}+{\mathbf{F}}_{12}{\mathbf{s}}_{Q}$ and ${\mathbf{F}}_{21}{\mathbf{s}}_{I}+{\mathbf{F}}_{22}{\mathbf{s}}_{Q}$. The output signal in complex form is thereby
${\mathbf{x}}={\mathbf{F}}_{11}{\mathbf{s}}_{I}+{\mathbf{F}}_{12}{\mathbf{s}}_{Q}+j({\mathbf{F}}_{21}{\mathbf{s}}_{I}+{\mathbf{F}}_{22}{\mathbf{s}}_{Q}).$ (23)
One can verify that the equivalent real-form signal is
$\bar{{\mathbf{x}}}=\left(\begin{array}[]{c}{\mathbf{x}}_{I}\\ {\mathbf{x}}_{Q}\end{array}\right)=\underbrace{\left(\begin{array}[]{cc}{\mathbf{F}}_{11}&{\mathbf{F}}_{12}\\ {\mathbf{F}}_{21}&{\mathbf{F}}_{22}\end{array}\right)}_{\bar{{\mathbf{F}}}}{\left(\begin{array}[]{c}{\mathbf{s}}_{I}\\ {\mathbf{s}}_{Q}\end{array}\right)},$ (30)
whose covariance matrix is $\bar{{\mathbf{Q}}}=\frac{1}{2}\bar{{\mathbf{F}}}\bar{{\mathbf{F}}}^{T}$. The matrix $\bar{{\mathbf{F}}}\in\mathbb{R}^{2N_{t}\times 2N_{s}}$ represents the “IQ-aware digital precoder”. By comparing (23) and (30) with their counterparts in (15) and (22), two important features of IQ-aware digital precoding can be observed. First, the complex signal ${\mathbf{x}}$ in (23) is no longer the product of a complex precoder and a complex symbol as in (15), making it possible for ${\mathbf{x}}$ to be general Gaussian distributed. Second, the equivalent real-valued precoder $\bar{{\mathbf{F}}}$ in (30), now is block-decoupled with independent block components, $\{{\mathbf{F}}_{11},{\mathbf{F}}_{12},{\mathbf{F}}_{21},{\mathbf{F}}_{22}\}$, which relaxes the block-coupled constraint of ${\mathbf{F}}_{r}$ in (22). These two features give the proposed IQ-aware digital precoding more flexibility in manipulating the transmitted signal ${\mathbf{x}}$, allowing for the utilization of SVD-based precoding technique to achieve the channel capacity.
@@
To be specific, we let $\bar{{\mathbf{V}}}\in\mathbb{R}^{2N_{t}\times 2N_{t}}$ be the right-singular matrix of $\bar{{\mathbf{H}}}$, and $\{\bar{\lambda}_{1},\cdots,\bar{\lambda}_{\min(N_{r},2N_{t})}\}$ the corresponding singular values in a descending order. The optimal IQ-aware digital precoder $\bar{{\mathbf{F}}}$ is represented as
@@
The power allocation matrix $\bar{{\mathbf{P}}}\overset{\Delta}{=}\mathrm{diag}\{\bar{p}_{1},\cdots,\bar{p}_{2N_{s}}\}$ is determined by the water-filling principle: $\bar{p}_{k}=\left(\mu-\frac{\sigma^{2}}{\bar{\lambda}_{k}^{2}}\right)^{+}$, where the water level $\mu$ is adjusted to satisfy the power constraint $\mathsf{Tr}\{\bar{{\mathbf{Q}}}\}=\frac{1}{2}\mathsf{Tr}\{\bar{{\mathbf{P}}}\}=\frac{1}{2}\sum_{k=1}^{2N_{s}}\bar{p}_{k}=P$. Consequently, by processing baseband data ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$ using $\{{\mathbf{F}}_{11},{\mathbf{F}}_{12},{\mathbf{F}}_{21},{\mathbf{F}}_{22}\}$, the channel capacity in (14) is achieved as:
$C=\sum_{k=1}^{2N_{s}}\frac{1}{2}\log\left(1+\frac{\bar{\lambda}_{k}^{2}\bar{p}_{k}}{\sigma^{2}}\right),$ (34)
where the number of real-data streams, $2N_{s}$, should not exceed the rank of the real-channel matrix, denoted as $\min(N_{r},2N_{t})$.
IV-B Channel Capacity Analysis
To gain more insights into atomic MIMO receivers, we analyze the asymptotic behavior of the channel capacity, $C$, at high SNRs with the consideration of the maximal number of multiplexed data streams, i.e., $2N_{s}=\min(N_{r},2N_{t})$.
Lemma 2.
Define $\text{SNR}=\frac{P}{\sigma^{2}}$. When $\text{SNR}\rightarrow+\infty$ and $2N_{s}=\min(N_{r},2N_{t})$, the channel capacity is asymptotically
$\boxed{C=\min\left(\frac{N_{r}}{2},N_{t}\right)\log\frac{\mathrm{SNR}}{\min\left(\frac{N_{r}}{2},N_{t}\right)}+\mathcal{O}(1).}$ (35)
Proof: Equation (35) holds because $p_{k}\rightarrow\frac{P}{N_{s}}$ when $\text{SNR}\rightarrow+\infty$ and $\log(1+x)\rightarrow\log(x)$ when $x\rightarrow+\infty$.  $\Box$
Lemma 2 suggests that the capacity of an atomic MIMO system is improved by $\min(N_{r}/2,N_{t})$ bit/s/Hz for every 3 dB increment in SNR, which corresponds to the DoFs of $\lim_{\text{SNR}\rightarrow\infty}\frac{C_{e}}{\log\text{SNR}}=\min\left(\frac{N_{r}}{2},N_{t}\right)$ [35]. This result reflects the fact that half DoFs are lost at the receiver side due to the absence of imaginary-part information in (8), while the DoFs remain intact at the transmitter side. Thereafter, a discount factor $\frac{1}{2}$ is imposed on $N_{r}$, giving the DoFs of $\min(\frac{N_{r}}{2},N_{t})$.
Figure 3: Comparison of DoFs between three cases: 1) $\min(N_{r},N_{t})$ for traditional MIMO, 2) $\min(\frac{N_{r}}{2},N_{t})$ for atomic MIMO, and 3) $\frac{1}{2}\min(N_{r},N_{t})$ for traditional MIMO with real symbols. The number of transimt antennas is set as $N_{t}=5$.
For a better understanding, we compare three DoFs, $\min(N_{r},N_{t})$, $\min(\frac{N_{r}}{2},N_{t})$, and $\frac{1}{2}\min(N_{r},N_{t})$, as a function of $N_{r}$ in Fig. 3. They correspond to the following three cases, respectively: 1) the traditional MIMO system, 2) the atomic MIMO system, 3) the In-phase transmission, where only real-valued symbols, such as ${\mathbf{s}}_{I}$, are transmitted through a traditional $N_{r}\times N_{t}$ MIMO channel. As $N_{r}$ increases, the DoFs $\min(\frac{N_{r}}{2},N_{t})$ gradually climb from $\frac{1}{2}\min(N_{r},N_{t})$ to $\min(N_{r},N_{t})$. Specifically, when $N_{r}<N_{t}$, we have $\min\left(\frac{N_{r}}{2},N_{t}\right)=\frac{1}{2}\min\left({N_{r}},N_{t}\right)=\frac{1}{2}N_{r}$. The DoFs of atomic MIMO systems are exactly half of that of conventional systems. In this context, it is sufficient to achieve the capacity of atomic receivers by transmitting $\frac{N_{r}}{2}$ complex-valued symbols, or equivalently $N_{r}$ real-valued symbols, such as ${\mathbf{s}}_{I}$. Moreover, as $N_{r}$ grows from $N_{t}$ to $2N_{t}$, the DoFs, $\min(N_{r},N_{t})$, saturate while the gap between $\min(N_{r},N_{t})$ and $\min(N_{r}/2,N_{t})$ narrows. Particularly, when $N_{r}\geq 2N_{t}$, we have $\min(\frac{N_{r}}{2},N_{t})=\min(N_{r},N_{t})=N_{t}$. This relationship indicates that an atomic receiver does not suffer any DoF loss compared to its traditional counterpart if the number of atomic antennas is at least twice the number of transmit antennas. This result is intuitive because $N_{r}$ real-valued observations are sufficient to recover $N_{t}\leq\frac{N_{r}}{2}$ complex-valued symbols. Consequently, under a same high SNR, the capacity gap between atomic and traditional MIMO systems tends to a constant independent of SNR. In summary, to fully exploit the multiplexing gain of an atomic MIMO system, it is necessary to have $N_{r}\geq 2N_{t}$.
@@ V IQ-Aware Hybrid Precoding
For massive MIMO systems employing large-scale antenna arrays, hybrid precoding that incorporates digital and analog precoders is a more practical solution than fully digital precoding[36, 37, 38, 39]. Its essential idea is to cascade a low-dimensional digital processor (i.e., digital precoder) and a high-dimensional analog PS network (i.e., analog precoder), so as to reduce the number of costly radio-frequency (RF) chains. In this section, we extend the fully digital IQ-aware precoding in the preceding section to realize IQ-aware hybrid precoding. Compared with traditional designs, a new challenge arises from the fact that the need of independent manipulation of IQ symbols cannot be implemented using analog circuits (i.e., a PS network).
To tackle this challenge, the proposed IQ-aware hybrid precoding architecture deploys a low-dimensional IQ-aware digital precoder to assist the analog PS network. As illustrated in Fig. 4, both the FC and SC phase-shifter networks are considered. Compared with the IQ-aware fully digital precoder in (30) where the entire precoding matrix is block-decoupled, the IQ-aware hybrid precoder is constructed by the product of a block-decoupled digital precoder and a block-coupled analog precoder. We will show that, despite the block-coupled constraint imposed on the analog part, the IQ-aware hybrid precoding is able to closely approximate the fully digital counterpart.
V-A IQ-Aware Hybrid Precoding with a FC-PS Network
V-A1 Problem statement
We first consider the IQ-aware hybrid precoding with a FC-PS network, where each RF chain is connected to all BS antennas via PSs. Denote the number of RF chains as $N_{\rm RF}$, satisfying $N_{s}\leq N_{\rm RF}\ll N_{t}$. In the baseband processor, an IQ-aware digital precoder $\bar{{\mathbf{D}}}$ of dimension $2N_{\rm RF}\times 2N_{s}$ is employed, with block components ${\mathbf{D}}_{k\ell}\in\mathbb{R}^{N_{\rm RF}\times N_{s}},k\in\{1,2\},\ell\in\{1,2\}$, to precode the symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$. The complex signal vector coming out of RF chains is thus ${\mathbf{s}}_{\rm rf}={\mathbf{D}}_{11}{\mathbf{s}}_{I}+{\mathbf{D}}_{12}{\mathbf{s}}_{Q}+j({\mathbf{D}}_{21}{\mathbf{s}}_{I}+{\mathbf{D}}_{22}{\mathbf{s}}_{Q})$. At the RF front-end, the analog precoder is expressed by a complex matrix ${\mathbf{A}}={\mathbf{A}}_{I}+j{\mathbf{A}}_{Q}\in\mathbb{C}^{N_{t}\times N_{\rm RF}}$. Given that PSs can only adjust the phases of signals, each element of ${\mathbf{A}}$ must satisfy the unit modulus constraint: $|{\mathbf{A}}_{m,n}|^{2}={\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\forall m,n$. Thereafter, the transmitted signal ${\mathbf{x}}$ is given as ${\mathbf{x}}={\mathbf{A}}{\mathbf{s}}_{\rm rf}$, whose real form is
The covariance matrix of $\bar{{\mathbf{x}}}$ is $\bar{{\mathbf{Q}}}=\frac{1}{2}\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\bar{{\mathbf{D}}}^{T}\bar{{\mathbf{A}}}^{T}$. Comparing the precoding models (30) and (42), one can observe that the analog part $\bar{{\mathbf{A}}}$ is subject to the block-coupled constraint, since analog PSs cannot distinguish the combined IQ signals. Fortunately, the hybrid precoder, $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}$, still has the ability to produce general Gaussian distributed signals $\bar{{\mathbf{x}}}$ because the digital part, $\bar{{\mathbf{D}}}$, is IQ-aware and has independent block components. This ability allows the accurate approximation of the IQ-aware fully digital precoder, $\bar{{\mathbf{F}}}$, by the hybrid counterpart, $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}$. To realize this objective, the joint optimization of the matrices $\bar{{\mathbf{A}}}$ and $\bar{{\mathbf{D}}}$ are formulated as
$\mathrm{s.t.}\>\>$ ${\mathbf{A}}_{I,m,n}^{2}+{\mathbf{A}}_{Q,m,n}^{2}=1,\>\forall m,n,$ (43a)
wherein (43b) is the total power constraint. In the subsequent discussions, we will devise an alternating minimization algorithm to solve $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}$ from problem (43) and use numerical simulations to show that the achieved rate of $\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}$ can tightly approach the channel capacity in (34).